Math 125 Quiz 2 (20 minutes)

Name  Student number  Quiz section:

There are three problems in this quiz (the other one is on the back of this page), total of 20 points. Write legibly and show your work fully.

1 (5 points) Find the derivative of f(x) = ∫_{x^{2}}^{e^{2x}}(ln t + sin t) dt.

2 (5 points) Evaluate the definite integral ∫0^{\frac{\sqrt{π}}{2}} x cos(x^{2}) dx.

3. Evaluate the indefinite integral.

(a) (5 points) ∫ x\sqrt{x + 1} dx;

(b) (5 points) ∫ \frac{(1 − x^{2})^{2}}{\sqrt{x}} dx.